OpenStudy (anonymous):

What is the GCF of 140 and 180? WILL MEDAL AND FAN.

OpenStudy (welshfella):

find the prime factors of the 2 numbers 140 = 2*2*5*7 180 = 2*2*5*9

OpenStudy (anonymous):

my options are 30, 20, 10, and 90

OpenStudy (welshfella):

the GCF is the numbers which are common to both factors here it is 2*2*5

OpenStudy (welshfella):

2* 2 * 5 = ?

OpenStudy (anonymous):

no what is the gcf 90, 30, 20, 10?

OpenStudy (anonymous):

20

OpenStudy (welshfella):

well as I said it is 2*2*5

OpenStudy (welshfella):

right

OpenStudy (welshfella):

do you follow the process OK?

OpenStudy (anonymous):

huh?

OpenStudy (welshfella):

do you see how i got 140 = 2*2*5*7 ?

OpenStudy (welshfella):

how ignorant!